Santiago Morgan is insured under a major medical expense insurance policy that includes a $200 calendar year deductible and a 20 percent coinsurance requirement from the insured. During the last year, Mr. Morgan incurred $1,000 in allowable medical expenses. Of this amount, Mr. Morgan is responsible for paying $640.00 $200.00 $160.00 $360.00
Answer #1

The deductible amount = 20% of $1,000 = $200

Amount after deductions = 1000 - 200 = $800

20% coinsurance requirement will require Santiago to also pay 20% of $800

So 20% of $800 = $160

So total amount that Mr. Morgan is responsible for paying

= 200 + 160

= $360